Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
nardecchia, palma - 20:33 Tuesday 23 January 2024 (63070) Print this report
CO2 actuators power change to compensate for 15W of input power

After having waited for the stabilization of NI/WI lasers power needed because of the replacement of the WI backup chiller, we increased the CO2 actuators power of about 25% wrt the previous values in order to follow the input power increase from 12W to 15W (+25%). This action allowed to reach the steady state at the beginning of the afternoon shift. 

 

At 10.45 UTC we started to change the powers. The previous and the current values are shown in the following table:


 

   

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.042 → 0.053      

0

1.38 → 1.73    

on the pickoff 

0.258 → 0.326       

0

0.225 → 0.282

N

on the ITF

0.050→ 0.063      

0.135+0.080=0.215 →0.17+0.080=0.250

0.86+1.80 =2.66 → 1.08+1.80=2.88 

on the pickoff 

0.308 →   0.39      

0.022+ 0.013 = 0.035→ 0.028+0.013 =0.041

0.14+0.293=0.433 → 0.176+0.293=0.470 

At 10.58 UTC all the flip mirrors were restored in the proper configuration to be able to start the lock acquisition after the INJ recovery.


 

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×